# Molecule Graph Networks with Many-body Equivariant Interactions

### v2

@@ Appendix A Mathematical Foundations and Analysis of the ENINet Architecture
Furthermore, by representing the linear transformations using their matrix forms, the matrix version of Equation (43) holds by replacing the composition operation $\circ$ with matrix multiplication $\cdot$ (cf. Definition S11).
@@
This section introduces key concepts from group representation theory that are essential for analyzing network architectures in this paper. Specifically, we focus on orthogonal groups, denoted ${\rm O}(n)$, where $n$ represents positive integers. The results presented here provide the foundation for examining the ${\rm O}(n)$-equivalence and related properties of the proposed ENINet architecture. Most of the foundational concepts and theories presented here can be found in standard textbooks on group representation theory, such as [66, 67, 68].
@@
Triples such as $(\mathbb{R},+,0)$, $(\mathbb{Z},+,0)$, and $(\mathbb{Q},+,0)$ are typical examples of commutative groups. In contrast, groups like $({\rm GL}(n),\circ,I_{n})$, $({\rm O}(n),\circ,I_{n})$, $({\rm SO}(n),\circ,I_{n})$, etc., are generally non-Abelian since matrix multiplication is generally non-commutative. In this paper, we focus on group representations of these non-commutative groups. Formulations of these non-Abelian groups are provided in the following examples (Examples S2 and S3).
@@
Let $(G,*,e)$ be a group. A non-empty subset $H$ of $G$ is called a subgroup of $G$ if $g^{-1}*h\in H$ whenever $g,h\in H$. In particular, the operation * forms a binary operation on $H$ with $e\in H$, such that the triple $(H,*,e)$ forms a group.
Example S3 provides some typical examples of subgroups of ${\rm GL}(n)$, such as the special linear group ${\rm SL}(n)$, the orthogonal group ${\rm O}(n)$, and the special orthogonal group ${\rm SO}(n)$, among others. Each of these groups corresponds to specific families of geometric transformations within the $n$-dimensional Euclidean space. Specifically, in this paper, we focus on the case $n=3$ for 3D molecular representations.
@@
  • (Orthogonal group) ${\rm O}(n)=\{A\in\mathbb{R}^{n\times n}\ |\ AA^{\top}=A^{\top}A=I_{n}\}$;
  • (Special orthogonal group) ${\rm SO}(n)=\{A\in{\rm O}(n)\ |\ \det(A)=1\}$.
In particular, ${\rm SL}(n)$, ${\rm O}(n)$, and ${\rm SO}(n)$ are subgroups of ${\rm GL}(n)$, and ${\rm SO}(n)$ is a subgroup of ${\rm O}(n)\cap{\rm SL}(n)$.
With various constraints on invertible matrices, the groups defined in Example S3 exhibit distinct geometric significance. For instance, by controlling the determinant of the matrices, elements of ${\rm SL}(n)$ preserve the “volumes” of objects within $\mathbb{R}^{n}$. On the other hand, matrices in the orthogonal group ${\rm O}(n)$ generalize the 3-dimensional rotation matrices in ${\rm O}(3)$. Specifically, matrices in ${\rm O}(n)$ are isometric linear transformations of $\mathbb{R}^{n}$ onto itself (Example S13).
@@
Every group homomorphism $f:G\rightarrow H$ between two groups $(G,*,e_{G})$ and $(H,\star,e_{H})$ sends the identity element $e_{G}$ in $G$ to the identity map $e_{H}$ in $H$, i.e., $f(e_{G})=e_{H}$. If $H$ is a subgroup of $G$, then the inclusion map $H\hookrightarrow G$ is a group homomorphism. In particular, the canonical group homomorphism ${\rm O}(n)\hookrightarrow{\rm GL}(n)$ is the primary focus of this paper.
@@
Following the discussion and conventions in Section A.3.1 and the above paragraph, any subgroup of ${\rm GL}(n)$ is naturally represented by the vector space $\mathbb{R}^{n}$ via the group homomorphism $G\hookrightarrow{\rm GL}(n)$. We formally state this fact by the following example.
@@
$\begin{bmatrix}\rho_{1}(g)&O&O&\cdots&O\\ O&\rho_{2}(g)&O&\cdots&O\\ O&O&\rho_{3}(g)&\cdots&O\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ O&O&O&\cdots&\rho_{m}(g)\\ \end{bmatrix},$
which is a block matrix consisting of $m^{2}$ linear transformations, where $O:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n_{j}}$ denote the zero matrices for $i,j\in\{1,2,...,m\}$.
@@
Because $\rho_{g}$ is invertible for every $g\in G$, the associated linear transformation $I_{d}\otimes\rho_{g}:\mathbb{R}^{d}\otimes\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{n}$ is also invertible. More precisely, the inverse map of $I_{d}\otimes\rho_{g}$ is exactly the tensor $I_{d}\otimes\rho_{g}^{-1}$, and this shows that $\varphi:G\rightarrow{\rm GL}(\mathbb{R}^{d}\otimes\mathbb{R}^{n})$ is well-defined. Furthermore, for $g,h\in G$, $\varphi(gh)=I_{d}\otimes\rho_{gh}=I_{d}\otimes(\rho_{g}\cdot\rho_{h})=(I_{d}\otimes\rho_{g})\cdot(I_{d}\otimes\rho_{h}))$, where $\cdot$ denotes the matrix product. Therefore, the map $\varphi:G\rightarrow{\rm GL}(\mathbb{R}^{d}\otimes\mathbb{R}^{n})$ is a group homomorphism. ∎
@@
In particular, if two group representations $\rho:G\rightarrow{\rm GL}(V)$ and $\varphi:G\rightarrow{\rm GL}(W)$ are equivalent, then the degree of $\rho$ and $\varphi$ are the same since there is an $\mathbb{R}$-isomorphism $T:V\rightarrow W$. In addition, if $\rho\sim\varphi$ and $\varphi\sim\sigma$ hold for representations $\rho:G\rightarrow{\rm GL}(V)$, $\varphi:G\rightarrow{\rm GL}(W)$, and $\sigma:G\rightarrow{\rm GL}(P)$, then $\rho\sim\sigma$.
@@
In what follows, we examine examples of equivariant functions that are fundamental to the analysis of the ENINet architecture. Specifically, we focus on ${\rm O}(n)$-equivariant functions, particularly in the case where $n=3$, and explore their connections to the ENINet architecture.
@@
Let $\rho:{\rm O}(n)\hookrightarrow{\rm GL}(n)$ be the canonical group representation of ${\rm O}(n)$. Then, $n\in{\rm{Hom}}_{\rho,\rho}(\mathbb{R}^{n},\mathbb{R}^{n})$.
@@
To prove that $\mathfrak{n}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is ${\rm O}(n)$-equivariant with respect to the representation $\rho$ and itself, one verifies that the diagram
@@
is commutative. First, if $\mathbf{x}=\mathbf{0}\in\mathbb{R}^{n}$, then $Q\cdot\mathfrak{n}(\mathbf{x})=Q\cdot\mathbf{0}=\mathbf{0}$ and $\mathfrak{n}(Q\cdot\mathbf{x})=\mathfrak{n}(Q\cdot\mathbf{0})=\mathfrak{n}(\mathbf{0})=\mathbf{0}$ since $Q$ is a linear transformation. Let $\mathbf{x}\in\mathbb{R}^{n}\setminus\{\mathbf{0}\}$. Then,
@@
for every $Q\in{\rm O}(n)$. One also notices that $Q\cdot\mathbf{x}\neq\mathbf{0}$ since $\mathbf{x}\neq\mathbf{0}$ and $Q$ is injective. Consequently, we conclude that the normalization function $\mathfrak{n}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is ${\rm O}(n)$-equivariant with respect to the representation $\rho$ and itself. ∎
@@
Let $G$ be a group and $\rho:{\rm O}(n)\rightarrow{\rm GL}(n)$ be a group representation. Let $d\in\mathbb{N}$ be a positive integer. Let $\varphi:{\rm O}(n)\rightarrow{\rm GL}(\mathbb{R}^{d}\otimes\mathbb{R}^{n})$ be the group representation defined in Definition S11. Let $f:\mathbb{R}^{d}\oplus\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{n}$ be the function that maps $(\mathbf{x},\mathbf{y})$ to the tensor $\mathbf{x}\otimes\mathbf{y}$. Then, the diagram
@@
is commutative for every matrix $Q\in{\rm O}(n)$. In particular, the function $f$ is ${\rm O}(n)$-equivariant with respect to the group representations $\pi_{d}\oplus\rho$ and $\varphi$, where $\pi_{d}:{\rm O}(n)\rightarrow{\rm GL}(\mathbb{R}^{d})$, $Q\mapsto I_{d}$ is the trivial representation (cf. Example S7).
@@
Equivariance describes a function $f:V\rightarrow W$ between representation spaces of a group $G$ that consistently transforms in accordance with the group representation acting on the domain elements. On the other hand, invariance focuses on whether the output of $f$ remains unchanged under transformations of the domain elements induced by the group. Actually, as shown in Proposition S2, by choosing an appropriate representation $\varphi$ within the space ${\rm{Hom}}_{\rho,\varphi}(V,W)$, invariant functions can be identified as a special case of equivariant functions.
@@
Let $G$ be a group, and let $\rho:G\rightarrow{\rm GL}(V)$ be a representation of $G$. A function $f:V\rightarrow W$ between real-valued vector spaces $V$ and $W$ is said to be G-invariant if $f=f\circ\rho_{g}$ for every $g\in G$.
@@
Let $G$ be a group, and let $\rho:G\rightarrow{\rm GL}(V)$ be a representation of $G$. A function $f:V\rightarrow W$ is $G$-invariant if and only if it is $G$-equivariant with respect to the representations $\rho$ and $\pi$, where $\pi:G\rightarrow{\rm GL}(W)$ is the trivial representation.
@@
If $f=f\circ\rho_{g}$ for every $g\in G$, then $\pi_{g}\circ f={\rm{id}}_{W}\circ f=f=f\circ\rho_{g}$ for every $g\in G$, and this shows that $f\in{\rm{Hom}}_{\rho,\pi}(V,W)$. Conversely, suppose $f\in{\rm{Hom}}_{\rho,\pi}(V,W)$, then $f={\rm{id}}_{W}\circ f=\pi_{g}\circ f=f\circ\rho_{g}$ for every $g\in G$, and this shows that $f$ is $G$-invariant. ∎
@@
The following are two typical examples of ${\rm O}(n)$-invariant functions between vector spaces: the inner product and norm functions on finite-dimensional $\mathbb{R}$-vector spaces, both of which are invariant under orthogonal transformations.
@@
Let $\rho:{\rm O}(n)\hookrightarrow{\rm GL}(n)$ be the canonical group representation of ${\rm O}(n)$, and let $\varphi:{\rm O}(n)\rightarrow{\rm GL}(1)$ be the trivial representation that sends each $Q$ to $1\in\mathbb{R}^{\times}={\rm GL}(1)$. Let $\|\cdot\|:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be norm function. Then, the diagram
@@
commutes for every orthogonal matrix $Q\in{\rm O}(n)$. In other words, the norm function $\|\cdot\|:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is ${\rm O}(n)$-equivariant with respect to the representations $\rho$ and $\varphi$.
@@
Similar to the proof of Example S12, one has $\varphi_{Q}(\|\mathbf{x}\|)=\|\mathbf{x}\|=\sqrt{\langle\mathbf{x},\mathbf{x}\rangle}$ and $\|Q\mathbf{x}\|=\sqrt{\langle Q\mathbf{x},Q\mathbf{x}\rangle}=\sqrt{\langle\mathbf{x},\mathbf{x}\rangle}=\|\mathbf{x}\|=\varphi_{Q}(\|\mathbf{x}\|)$. Therefore, $\|\cdot\|:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is ${\rm O}(n)$-equivariant with respect to $\rho$ and $\varphi$. ∎
@@
Let $G$ be a group and $\rho:G\rightarrow{\rm GL}(V),\varphi:G\rightarrow{\rm GL}(W)$ be representations. Then ${\rm{Hom}}_{\rho,\varphi}(V,W)$ forms a vector space over $\mathbb{R}$, where $(f_{1}+f_{2})(v):=f_{1}(v)+f_{2}(v)$ and $(rf)(v):=rf(v)$ whenever $v\in V$, $f,f_{1},f_{2}\in{\rm{Hom}}_{\rho,\varphi}(V,W)$, and $r\in\mathbb{R}$.
@@
Let $0:V\rightarrow W$ be the zero function that maps every $v\in V$ to the zero element in $W$. Then for every $v\in V$ and $g\in G$, $(0\circ\rho_{g})(0)=0=\varphi_{g}(0)=(\varphi_{g}\circ 0)(v)$ since $\varphi_{g}$ is a linear transformation. This shows that $0\in{\rm{Hom}}_{\rho,\varphi}(V,W)$, where $00+f=f=f+$ 0 whenever $f\in{\rm{Hom}}_{\rho,\varphi}(V,W)$. Because $f_{1},f_{2}\in{\rm{Hom}}_{\rho,\varphi}(V,W)$ and $\rho_{g},\varphi_{g}$ are $\mathbb{R}$-linear, $(\varphi_{g}\circ(f_{1}+f_{2}))(v)=\varphi_{g}(f_{1}(v)+f_{2}(v))=\varphi_{g}(f_{1}(v))+\varphi_{g}(f_{2}(v))=f_{1}(\rho_{g}(v))+f_{2}(\rho_{g}(v))=(f_{1}+f_{2})(\rho_{g}(v))$ whenever $v\in V$. This shows that $f_{1}+f_{2}\in{\rm{Hom}}_{\rho,\varphi}(V,W)$. On the other hand, $rf\in{\rm{Hom}}_{\rho,\varphi}(V,W)$ since $(\varphi_{g}\circ(rf))(v)=r\cdot\varphi_{g}(f(v))=r\cdot f(\rho_{g}(v))=((rf)\circ\rho_{g})(v)$ whenever $v\in V$. Therefore, equipped with the addition and scalar multiplication operations, ${\rm{Hom}}_{\rho,\varphi}(V,W)$ forms a vector space over $\mathbb{R}$. ∎
@@
Let $G$ be a group, and let $\rho:G\rightarrow{\rm GL}(V)$, $\varphi:G\rightarrow{\rm GL}(W)$, and $\psi:G\rightarrow{\rm GL}(P)$ be representations. If $f:V\rightarrow W$ and $g:W\rightarrow P$ are $G$-equivariant, then $g\circ f$ is also $G$-equivariant. In other words, $g\circ f\in{\rm{Hom}}_{\rho,\psi}(V,P)$ whenever $f\in{\rm{Hom}}_{\rho,\varphi}(V,W)$ and $g\in{\rm{Hom}}_{\varphi,\psi}(W,P)$.
@@
Based on the graph model settings of the input molecule, we can consider the feature representations of graph edges and investigate the ${\rm O}(3)$-equivariance of feature aggregation. Formally, let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a finite, undirected, simple graph, where $\mathcal{V}=\{1,2,\dots,n\}$ consists of $n$ vertices, and each vertex is assigned a coordinate $\mathbf{r}_{i}=(x_{i},y_{i},z_{i})\in\mathbb{R}^{3}$, satisfying $\mathbf{r}_{i}\neq\mathbf{r}_{j}$ whenever $i\neq j$. Furthermore, for distinct vertices $i$ and $j$, we denote $\vec{\textbf{r}}_{ji}:=(x_{ji},y_{ji},z_{ji})=\mathbf{r}_{j}-\mathbf{r}_{i}$, which records the directional information from $\mathbf{r}_{j}$ to $\mathbf{r}_{i}$. For every fixed $i\in\mathcal{V}$ and $j\in\mathcal{V}$ that satisfies $\{i,j\}\in\mathcal{E}$, the directional feature of the edge $\{i,j\}$ is defined as
@@
where $\vec{\mathbf{f}}_{ij}\in\mathbb{R}^{d}$ is a $d$-dimensional feature vector of the directional edge from $j$ to $i$. In particular, for a fixed vertex $i\in\mathcal{V}$ and a given $j\in\mathcal{N}(i)$, the information in (44) can be encoded by the function

### v3

@@ Appendix A Mathematical Foundations and Analysis of the ENINet Architecture
Furthermore, by representing the linear transformations via their matrix forms with respect to consistent bases (e.g., the standard bases in Euclidean spaces), the matrix version of Equation (45) holds, where the composition operation $\circ$ is replaced by matrix multiplication $\cdot$ (cf. Definition S11).
@@
This section introduces key concepts from group representation theory that are essential for analyzing network architectures in this paper. Specifically, we focus on orthogonal groups, denoted $\operatorname{O}(n)$, where $n$ represents positive integers. The results presented here provide the foundation for examining the $\operatorname{O}(n)$-equivariance and related properties of the proposed ENINet architecture. Most of the foundational concepts and theories presented here can be found in standard textbooks on group representation theory, such as [68, 69, 70].
@@
Triples such as $(\mathbb{R},+,0)$, $(\mathbb{Z},+,0)$, and $(\mathbb{Q},+,0)$ are typical examples of commutative groups. In contrast, groups like $({\rm GL}(n),\circ,I_{n})$, $(\operatorname{O}(n),\circ,I_{n})$, $({\rm SO}(n),\circ,I_{n})$, etc., are generally non-Abelian since matrix multiplication is generally non-commutative. In this paper, we focus on group representations of these non-commutative groups. Formulations of these non-Abelian groups are provided in the following examples (Examples S2 and S3).
@@
[…]
Example S3 provides some typical examples of subgroups of ${\rm GL}(n)$, such as the special linear group ${\rm SL}(n)$, the orthogonal group $\operatorname{O}(n)$, and the special orthogonal group ${\rm SO}(n)$. Each of these groups corresponds to specific families of geometric transformations within the $n$-dimensional Euclidean space. Specifically, in this paper, we focus on the case $n=3$ for representations of 3D molecular structures.
@@
  • (Orthogonal group) $\operatorname{O}(n)=\{A\in\mathbb{R}^{n\times n}\ |\ AA^{\top}=A^{\top}A=I_{n}\}$;
  • (Special orthogonal group) ${\rm SO}(n)=\{A\in\operatorname{O}(n)\ |\ \det(A)=1\}$.
In particular, ${\rm SL}(n)$, $\operatorname{O}(n)$, and ${\rm SO}(n)$ are subgroups of ${\rm GL}(n)$, and ${\rm SO}(n)$ is a subgroup of $\operatorname{O}(n)\cap{\rm SL}(n)$.
With various constraints on invertible matrices, the groups defined in Example S3 exhibit distinct geometric significance. For instance, by controlling the determinant of the matrices, elements of ${\rm SL}(n)$ preserve the “volumes” of objects within $\mathbb{R}^{n}$. On the other hand, matrices in the orthogonal group $\operatorname{O}(n)$ generalize the 3-dimensional rotation matrices in $\operatorname{O}(3)$. Specifically, matrices in $\operatorname{O}(n)$ are isometric linear transformations of $\mathbb{R}^{n}$ onto itself (Example S13).
@@
Every group homomorphism $f:G\rightarrow H$ between two groups $(G,*,e_{G})$ and $(H,\star,e_{H})$ sends the identity element $e_{G}$ in $G$ to the identity map $e_{H}$ in $H$, i.e., $f(e_{G})=e_{H}$. If $H$ is a subgroup of $G$, then the inclusion map $H\hookrightarrow G$ is a group homomorphism. In particular, the canonical group homomorphism $\operatorname{O}(n)\hookrightarrow{\rm GL}(n)$ is the primary focus of this paper.
@@
Following the discussion and conventions in Section A.3.1 and the above paragraph, any subgroup $G$ of ${\rm GL}(n)$ is naturally represented by the vector space $\mathbb{R}^{n}$ via the group homomorphism $G\hookrightarrow{\rm GL}(n)$. We formally state this fact by the following example.
@@
$\begin{bmatrix}\rho_{1}(g)&O_{12}&O_{13}&\cdots&O_{1m}\\ O_{21}&\rho_{2}(g)&O_{23}&\cdots&O_{2m}\\ O_{31}&O_{32}&\rho_{3}(g)&\cdots&O_{3m}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ O_{m1}&O_{m2}&O_{m3}&\cdots&\rho_{m}(g)\\ \end{bmatrix},$
which is a block matrix consisting of $m^{2}$ linear transformations, where matrices $O_{ji}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n_{j}}$ denote the zero matrices for $i,j\in\{1,2,...,m\}$.
@@
Because $\rho_{g}$ is invertible for every $g\in G$, the associated linear transformation $I_{d}\otimes\rho_{g}:\mathbb{R}^{d}\otimes\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{n}$ is also invertible. More precisely, the inverse map of $I_{d}\otimes\rho_{g}$ is exactly the tensor $I_{d}\otimes\rho_{g}^{-1}$, and this shows that $\varphi:G\rightarrow{\rm GL}(\mathbb{R}^{d}\otimes\mathbb{R}^{n})$ is well-defined. Furthermore, for $g,h\in G$, $\varphi(gh)=I_{d}\otimes\rho_{gh}=I_{d}\otimes(\rho_{g}\cdot\rho_{h})=(I_{d}\otimes\rho_{g})\cdot(I_{d}\otimes\rho_{h})$, where $\cdot$ denotes the matrix product. Therefore, the map $\varphi:G\rightarrow{\rm GL}(\mathbb{R}^{d}\otimes\mathbb{R}^{n})$ is a group homomorphism. ∎
@@
In particular, if two group representations $\rho:G\rightarrow{\rm GL}(V)$ and $\varphi:G\rightarrow{\rm GL}(W)$ are equivalent, then the degrees of $\rho$ and $\varphi$ are the same since there is an $\mathbb{R}$-isomorphism $T:V\rightarrow W$. In addition, if $\rho\sim\varphi$ and $\varphi\sim\sigma$ hold for representations $\rho:G\rightarrow{\rm GL}(V)$, $\varphi:G\rightarrow{\rm GL}(W)$, and $\sigma:G\rightarrow{\rm GL}(P)$, then $\rho\sim\sigma$.
@@
In what follows, we examine examples of equivariant functions that are fundamental to the analysis of the ENINet architecture. Specifically, we focus on $\operatorname{O}(n)$-equivariant functions, particularly in the case where $n=3$, and explore their connections to the ENINet architecture.
@@
Let $\rho:\operatorname{O}(n)\hookrightarrow{\rm GL}(n)$ be the canonical group representation of $\operatorname{O}(n)$. Then, $\mathfrak{n}\in{\rm{Hom}}_{\rho,\rho}(\mathbb{R}^{n},\mathbb{R}^{n})$.
@@
To prove that $\mathfrak{n}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is $\operatorname{O}(n)$-equivariant with respect to the representation $\rho$ and itself, one verifies that the diagram
@@
is commutative for every $Q\in\operatorname{O}(n)$.. First, if $\mathbf{x}=\mathbf{0}\in\mathbb{R}^{n}$, then $Q\cdot\mathfrak{n}(\mathbf{x})=Q\cdot\mathbf{0}=\mathbf{0}$ and $\mathfrak{n}(Q\cdot\mathbf{x})=\mathfrak{n}(Q\cdot\mathbf{0})=\mathfrak{n}(\mathbf{0})=\mathbf{0}$ since $Q$ is a linear transformation. Let $\mathbf{x}\in\mathbb{R}^{n}\setminus\{\mathbf{0}\}$. Then,
@@
for every $Q\in\operatorname{O}(n)$. One also notices that $Q\cdot\mathbf{x}\neq\mathbf{0}$ since $\mathbf{x}\neq\mathbf{0}$ and $Q$ is injective. Consequently, we conclude that the normalization function $\mathfrak{n}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is $\operatorname{O}(n)$-equivariant with respect to the representation $\rho$ and itself. ∎
@@
Let $G$ be a group and $\rho:\operatorname{O}(n)\rightarrow{\rm GL}(n)$ be a group representation. Let $d\in\mathbb{N}$ be a positive integer. Let $\varphi:\operatorname{O}(n)\rightarrow{\rm GL}(\mathbb{R}^{d}\otimes\mathbb{R}^{n})$ be the group representation defined in Definition S11. Let $f:\mathbb{R}^{d}\oplus\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{n}$ be the function that maps $(\mathbf{x},\mathbf{y})$ to the tensor $\mathbf{x}\otimes\mathbf{y}$. Then, the diagram
@@
is commutative for every matrix $Q\in\operatorname{O}(n)$. In particular, the function $f$ is $\operatorname{O}(n)$-equivariant with respect to the group representations $\pi_{d}\oplus\rho$ and $\varphi$, where $\pi_{d}:\operatorname{O}(n)\rightarrow{\rm GL}(\mathbb{R}^{d})$, $Q\mapsto I_{d}$ is the trivial representation (cf. Example S7).
@@
Equivariance describes a function $f:V\rightarrow W$ between representation spaces of a group $G$ that consistently transforms in accordance with the group representations acting on the domain elements. On the other hand, invariance focuses on whether the output of $f$ remains unchanged under transformations of the domain elements induced by the group. Actually, as shown in Proposition S2, by choosing an appropriate representation $\varphi$ within the space ${\rm{Hom}}_{\rho,\varphi}(V,W)$, invariant functions can be identified as a special case of equivariant functions.
@@
Let $G$ be a group, and let $\rho:G\rightarrow{\rm GL}(V)$ be a representation of $G$. A function $f:V\rightarrow W$ between real-valued vector spaces $V$ and $W$ is said to be G-invariant with respect to the representation $\rho$ if $f=f\circ\rho_{g}$ for every $g\in G$.
@@
Let $G$ be a group, and let $\rho:G\rightarrow{\rm GL}(V)$ be a representation of $G$. A function $f:V\rightarrow W$ is $G$-invariant with respect to $\rho$ if and only if it is $G$-equivariant with respect to the representations $\rho$ and $\pi$, where $\pi:G\rightarrow{\rm GL}(W)$ is the trivial representation.
@@
If $f=f\circ\rho_{g}$ for every $g\in G$, then $\pi_{g}\circ f={\rm{id}}_{W}\circ f=f=f\circ\rho_{g}$ for every $g\in G$, and this shows that $f\in{\rm{Hom}}_{\rho,\pi}(V,W)$. Conversely, suppose $f\in{\rm{Hom}}_{\rho,\pi}(V,W)$, then $f={\rm{id}}_{W}\circ f=\pi_{g}\circ f=f\circ\rho_{g}$ for every $g\in G$, and this shows that $f$ is $G$-invariant with respect to the representation $\rho$. ∎
@@
The following are two typical examples of $\operatorname{O}(n)$-invariant functions between vector spaces: the inner product and norm functions on finite-dimensional $\mathbb{R}$-vector spaces, both of which are invariant under orthogonal transformations.
@@
Let $\rho:\operatorname{O}(n)\hookrightarrow{\rm GL}(n)$ be the canonical group representation of $\operatorname{O}(n)$, and let $\varphi:\operatorname{O}(n)\rightarrow{\rm GL}(1)$ be the trivial representation that sends each $Q$ to $1\in\mathbb{R}^{\times}={\rm GL}(1)$. Let $\|\cdot\|:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be norm function. Then, the diagram
@@
commutes for every orthogonal matrix $Q\in\operatorname{O}(n)$. In other words, the norm function $\|\cdot\|:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is $\operatorname{O}(n)$-equivariant with respect to the representations $\rho$ and $\varphi$, i.e., the norm function $\|\cdot\|:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is $\operatorname{O}(n)$-invariant with respect to the representation $\rho$.
@@
Similar to the proof of Example S12, one has $\varphi_{Q}(\|\mathbf{x}\|)=\|\mathbf{x}\|=\sqrt{\langle\mathbf{x},\mathbf{x}\rangle}$ and $\|Q\mathbf{x}\|=\sqrt{\langle Q\mathbf{x},Q\mathbf{x}\rangle}=\sqrt{\langle\mathbf{x},\mathbf{x}\rangle}=\|\mathbf{x}\|=\varphi_{Q}(\|\mathbf{x}\|)$. Therefore, $\|\cdot\|:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is $\operatorname{O}(n)$-equivariant with respect to $\rho$ and $\varphi$. ∎
@@
[…]
@@
[…]
@@
Let $G$ be a group, and let $\rho:G\rightarrow{\rm GL}(V)$, $\varphi:G\rightarrow{\rm GL}(W)$, and $\psi:G\rightarrow{\rm GL}(P)$ be representations. If $f\in{\rm{Hom}}_{\rho,\varphi}(V,W)$ and $g\in{\rm{Hom}}_{\varphi,\psi}(W,P)$, then $g\circ f\in{\rm{Hom}}_{\rho,\psi}(V,P)$.
@@
Based on the graph model settings of the input molecule, we can consider the feature representations of graph edges and investigate the $\operatorname{O}(3)$-equivariance of feature aggregation. Formally, let $\mathscr{G}=(\mathscr{V},\mathscr{E})$ be a finite, undirected, simple graph, where $\mathscr{V}=\{1,2,\dots,n\}$ consists of $n$ vertices, and each vertex is assigned a coordinate $\mathbf{r}_{i}=(x_{i},y_{i},z_{i})\in\mathbb{R}^{3}$, satisfying $\mathbf{r}_{i}\neq\mathbf{r}_{j}$ whenever $i\neq j$. Furthermore, for distinct vertices $i$ and $j$, we denote $\vec{\textbf{r}}_{ji}:=(x_{ji},y_{ji},z_{ji})=\mathbf{r}_{i}-\mathbf{r}_{j}$, which records the directional information from $\mathbf{r}_{j}$ to $\mathbf{r}_{i}$. For every fixed $i\in\mathscr{V}$ and $j\in\mathscr{V}$ that satisfies $\{i,j\}\in\mathscr{E}$, the directional feature of the edge $\{i,j\}$ is defined as
@@
where $\vec{\mathbf{f}}_{ji}\in\mathbb{R}^{d}$ is a $d$-dimensional feature vector of the directional edge from $j$ to $i$. In particular, for a fixed vertex $i\in\mathscr{V}$ and a given $j\in\mathscr{N}(i)$, the information in (46) can be encoded by the function